OpenStudy (ttop0816):

Which of the following is a solution of z^6 = 64i? 2(cos105° + isin105°) 2(cos120° + isin120°) 2(cos135° + isin135°) 8(cos15° + isin15°)

hartnn (hartnn):

can you write "i" in the form of \(\cos \theta +i\sin \theta \) what value of theta will u take ?

hartnn (hartnn):

\(z^6 = 64(\cos \theta +i\sin \theta ) \\ z =2 (cos (\theta/6) + i \sin (\theta/6) )\) using DeMovire's Theorem, You only need theta :)

hartnn (hartnn):

Divide the theta by 6, and you will get your angle :)

OpenStudy (ttop0816):

THANK YOU!!(:

hartnn (hartnn):

welcome ^_^